OpenStudy (surana):

For the parent function y /- positive x, what effect does a value of b = -2 have on the graph?

OpenStudy (surana):

Technically, it's \[y = \sqrt{x}\]

OpenStudy (surana):

Possible answers are... A: Horizontal compression by a factor of 1/2 and a reflection across the y-axis. B: Horizontal compression by a factor of 1/2 and a reflection across the x-axis. C: Horizontal compression by a factor of -1/2 and a reflection across the y-axis. D: Horizontal compression by a factor of -1/2 and a reflection across the x-axis.

OpenStudy (rational):

\[y = a\sqrt{\color{blue}{b}(x-h)}+k\] is that the \(\color{blue}{b}\) you talking about ?

OpenStudy (surana):

I think that's how it's supposed to go, radicals aren't exactly something I'm good at working with.

OpenStudy (rational):

then \(b=-2\) copresses the parent graph horizontally by a factor of 1/2, then reflects across y axis

OpenStudy (surana):

So what you mean is that it ends up going across the graph?

OpenStudy (rational):

look at ur options

OpenStudy (surana):

So you're saying B?

OpenStudy (rational):

A

OpenStudy (surana):

Oh. Yeah, you're right. I had to retake a look at it. Thank you!

OpenStudy (rational):

yw it makes more sense if you graph \(y=\sqrt{x}\) and \(y = \sqrt{-2x}\) and see how they differ